OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10) A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6

jimthompson5910 (jim_thompson5910):

y = x + 8 has a slope of 1

jimthompson5910 (jim_thompson5910):

So any parallel line also has a slope of 1

jimthompson5910 (jim_thompson5910):

y = mx+b y = 1x+b 10 = 1(4) + b 10 = 4 + b Now solve for b

OpenStudy (anonymous):

@jim_thompson5910 - B= 6

jimthompson5910 (jim_thompson5910):

so the equation is y = 1x + 6 or just y = x+6

OpenStudy (anonymous):

@jim_thompson5910 Do you see a difference between B and D

jimthompson5910 (jim_thompson5910):

no, B and D look the same...hmm

jimthompson5910 (jim_thompson5910):

typo maybe? or maybe something is missing

OpenStudy (anonymous):

Thank You!!!

jimthompson5910 (jim_thompson5910):

yw, so was there a typo?

OpenStudy (anonymous):

No, That is what it says on my paper. So If i get it wrong I will tell m teacher he missed typed.

jimthompson5910 (jim_thompson5910):

i gotcha, that's what I would do too

OpenStudy (anonymous):

:) thanks for your help